# Vector-Quantized Vision Foundation Models for Object-Centric Learning

### v5

@@ 1. Introduction
Objects can form highly diverse visual scenes through arrangements and combinations. But mainstream methods based on feature patches or a single feature vector disregard such compositionality. Inspired by human vision cognition (Bar, 2004; Cavanagh, 2011; Palmeri and Gauthier, 2004), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects. Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a), as well as in interactions between visual modality and other modalities (Zhang et al., 2024; Wang et al., 2024).
Existing OCL methods typically adopt an encoder-aggregator-decoder architecture (Locatello et al., 2020). Firstly, the encoder transforms input image or video frame pixels into a dense feature map. Then, the aggregator sparsifies this feature map into feature vectors via Slot Attention (Locatello et al., 2020) with initial slots as the query. Lastly, the decoder reconstructs the input in some form from these aggregated slots, to provide the self-supervised training signal.
OCL relies on pixel textures to discover objects. The early milestone (Locatello et al., 2020) reconstructs input pixels as supervision, usually failing on realistic objects. Some (Kipf et al., 2022; Elsayed et al., 2022) reconstruct optical flow or depth map to mitigate textural noises, at the cost of expensive annotations. Some (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, whose super-pixels are codebook codes, thus suppressing pixel redundancy and facilitating aggregation from features into slots. Recent advances (Seitzer et al., 2023; Wu et al., 2023b) use Vision Foundation Models (VFMs) (Caron et al., 2021; Oquab et al., 2023) to extract features with better object separability, boosting OCL significantly.
@@ 2. Related Work
OCL encoding. Early milestone methods like IODINE(Greff et al., 2019) and SA (Locatello et al., 2020) use small naive CNNs (Krizhevsky et al., 2012) as OCL encoder. Followups like SAVi (Kipf et al., 2022), SAVi++ (Elsayed et al., 2022), SLATE (Singh et al., 2022a) and STEVE (Singh et al., 2022c) employ pretrained ResNets (He et al., 2016), and fine-tune them on OCL datasets. State-of-the-arts like SlotDiffusion (Wu et al., 2023b) and DINOSAUR (Seitzer et al., 2023) utilize VFMs like DINO (Caron et al., 2021) and DINO2 (Oquab et al., 2023) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly. SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also well recognized VFMs yet remain unexploited in OCL setting. Our VVO supports various VFMs for OCL encoding.
OCL aggregation. SlotAttention (Locatello et al., 2020) is the footstone for mainstream OCL methods. Subsequent works like BO-QSA (Jia et al., 2023), ISA (Biza et al., 2023) and SysBind (Singh et al., 2022b) are all its variants, which are designed without changing the external interface. But considering their performance boosts, we only integrate BO-QSA by default.
OCL decoding. With SlotAttention as the aggregator, the decoder and its reconstruction target affect OCL performance the most, as it is the source of supervision. Mixture-based decoding, used in SAVi, SAVi++, DINOSAUR and VideoSAUR (Zadaianchuk et al., 2024), decodes each slot’s spatial broadcast (Watters et al., 2019) using naive CNNs or MLPs, and mixes them with corresponding weights into the reconstruction. Transformer-based decoding, used in SLATE, STEVE and SPOT (Kakogeorgiou et al., 2024), reconstructs VAE representation of the input auto-regressively with slots as the condition. Diffusion-based decoding in LSD (Jiang et al., 2023) and SlotDiffusion drives slots to recover noise added to the input’s VAE representation. Our VVO supports all these types of OCL decoding.
VAE for OCL. Variational Autoencoders (VAEs), like dVAE (Im Im et al., 2017) in SLATE and VQ-VAE (Van Den Oord et al., 2017) in SlotDiffusion, are employed to produce reconstruction targets for OCL training. Since these VAEs are designed for image generation, some methods adapt them for OCL. Inspired by channel or weight grouping, GDR (Zhao et al., 2024a) decomposes features into attributes and combine them to produce VAE representation as reconstruction targets to guide OCL better. MSF (Zhao et al., 2024b) firstly exploits the multi-scale idea in the OCL setting with VAE-specific designs. Based on recent advancement RVQ (Yang et al., 2023) and SimVQ (Zhu et al., 2024), we design our own VQ variant for OCL.
@@ 3.1. Unify OCL
where $\bm{\phi}_{\mathrm{e}}$ can be parameterized as pretrained VFMs, like DINO (Caron et al., 2021), DINO2 (Oquab et al., 2023), SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024). As OCL relies on textures to separate objects, $\bm{\phi}_{\mathrm{e}}$ should handle complex textures of objects, making VFMs necessary here. We will explain it in Section 3.2.
@@
where $\bm{\phi}_{\mathrm{a}}$ can be parameterized as widely adopted SlotAttention (Locatello et al., 2020) and its variants, which is some cross attention with $\bm{S}_{0}$ as queries and $\bm{Z}$ as keys and values. $\bm{M}$ is the binarized attention map thus intuitively reflects how well objects are represented by slots.
For video OCL, there is a recurrent module transitioning current slots $\bm{S}$ into new queries $\bm{S}_{0}$ for the next time step. Such module can be parameterized as a Transformer encoder block (Vaswani et al., 2017).
@@
where $\bm{\phi}_{\mathrm{q}}$ can be parameterized as some Vector Quantization (VQ) (Im Im et al., 2017; Van Den Oord et al., 2017). But we meticulously design our own VQ variant, as detailed in Section 3.2. $\bm{\phi}_{\mathrm{q}}$ is pretrained in VAE framework and is frozen afterwards, where the encoder is shared from the frozen OCL encoder $\bm{\phi}_{\mathrm{e}}$, and the decoder is a typical VAE decoder. We will explain why not use a separate typical VAE encoder in Section 3.2.
@@
Here $\bm{\phi}_{\mathrm{d}}$ can be parameterized as (i) a CNN or MLP for mixture decoding (Kipf et al., 2022; Seitzer et al., 2023), where $\bm{Q}$ is destructed to height and width, and $\bm{S}$ is spatially broadcast (Watters et al., 2019) into this shape and then decoded into components being mixed together; (ii) or a Transformer decoder for auto-regressive decoding (Singh et al., 2022a; Kakogeorgiou et al., 2024), where $\bm{Q}$ is destructed with causal masking as the query and $\bm{S}$ is the key and value; (iii) or a conditional Diffusion model for diffusion decoding (Wu et al., 2023b; Jiang et al., 2023), where $\bm{Q}$ is destructed with noise as the input and $\bm{S}$ is the condition.
Reconstructing $\bm{Q}$ using $\bm{S}$ drives $\bm{S}$ to aggregate as much object information as possible. Thus, a good reconstruction target, like in (Zhao et al., 2024a, b), is very important. We will explain this in Section 3.2.
@@ 3.2. Utilize VFMs in OCL
We directly extract the feature map $\bm{Z}$ from the input $\bm{X}$ using VFMs as $\bm{\phi}_{\mathrm{e}}$, where DINO2 ViT and SAM2 encoder are chosen and experimented thoroughly. No extra position encoding is needed here like in SA (Locatello et al., 2020) because these VFMs already contain the positional information required in $\bm{\phi}_{\mathrm{a}}$. Since $\bm{\phi}_{\mathrm{e}}$ is frozen, We further use a trainable linear layer to adjust $\bm{Z}$ slightly.
As shown in Figure 1, VFM representations have better objectness than non-VFMs, even under naive kMeans clustering [^1] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html. OCL aggregation is essentially a clustering whose initial centroids are trainable (Jia et al., 2023). Thus, we expect $\bm{\phi}_{\mathrm{a}}$ to aggregate VFM’s $\bm{Z}$ into slots $\bm{S}$ better under queries $\bm{S}_{0}$. Previous methods like DINOSAUR (Seitzer et al., 2023) have already exploited this but without any reason.
@@
Our VQ’s codebook follows SimVQ (Zhu et al., 2024). We predefine $m=4096$ template features, i.e., a codebook $\bm{T}_{0}\in\mathbb{R}^{m\times c_{0}}$, which are randomly initialized and remain frozen. In vector quantization, we project $\bm{T}$ with a pre-trainable linear layer and match it with the adjusted $\bm{Z}$:
@@
where $\mathrm{softmax}(\cdot)$ is calculated along the channel dimension; $\bm{P}$ is the match probabilities; $\bm{I}\in\mathbb{R}^{h\times w}$ is the matched code indexes; $\mathrm{argmax}(\cdot)$ is calculated along the channel dimension; $\mathrm{index}(\cdot,\cdot)$ is operated along code number dimension. The typical STE (Bengio et al., 2013) on $\bm{Q}$, needed in pre-training, can be skipped during OCL training.
@@
where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is Gumbel noise and $\tau$ is the temperature. Training-time annealing residual connection (Zhao et al., 2024a, b) is added after Equation 10 to stabilize pre-training:
@@
where $\alpha$ is scheduled from 1 to 0 during pre-training using cosine-annealing. Besides typical losses of reconstruction, alignment and commitment (Van Den Oord et al., 2017), we regularize the adjusted $\bm{Z}$ to be normal:
@@
With all samples’ feature maps $\bm{Z}$ being represented with one codebook $\bm{T}$, the quantization $\bm{Q}$ naturally gains cross-sample consistency, helping the aggregation with queries $\bm{S}_{0}$, which are also shared across samples. Such tokenization is compatible with both regression and classification decoding. In contrast, methods like SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b) are faced with distribution gaps between $\bm{Q}$ and $\bm{Z}$, shown in Figure 1, due to separate VAE and OCL encoders. Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen OCL supervision.
@@ 3.3. Compare Architectures
- SLATE0 (the official version) (Singh et al., 2022a): (1) VFM features are NOT directly fed to the aggregator. (2) VFM features are shared discretized to scalar numbers and re-embedded into features to be learned latter. This loses much information of VFM features.
- SLATE (the improved version; adopted here) / STEVE (Jia et al., 2023; Singh et al., 2022c; Wu et al., 2023b): (1) Same as VVO. (2) Reconstruction targets are discretized from features of separate VAE encoding, not quantized from features shared from OCL encoding, causing optimization noises.
- SlotDiffusion (Wu et al., 2023b) / LSD (Jiang et al., 2023): (1) Same as VVO. (2) Reconstruction targets are quantized from separate VAE encoding features, not sharing OCL encoding features, causing optimization noises.
- DINOSAUR (Seitzer et al., 2023) / VideoSAUR (Zadaianchuk et al., 2024) / SPOT (Kakogeorgiou et al., 2024): (1) Same as VVO. (2) Reconstruction targets are shared from VFM features of OCL encoding without quantization, causing optimization noises.
@@
- Mixture-based OCL decoders, e.g., CNN (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), MLP (Seitzer et al., 2023) and SlotMixer (Zadaianchuk et al., 2024), are originally designed for continuous features, thus are compatible with our shared quantized VFM features.
- Auto-regressive OCL decoders, e.g., the Transformer decoder (Singh et al., 2022a, c) and Transformer9 (Kakogeorgiou et al., 2024), are designed for discretized features, while also showing applicability to continuous features (Seitzer et al., 2023; Kakogeorgiou et al., 2024). Thus they are applicable to our shared quantized VFM features.
- Diffusion-based OCl decoders, e.g., conditional Diffusion (Wu et al., 2023b; Jiang et al., 2023), work on low-dimension features, necessitating our shared quantization on the continuous high-dimensional VFMs features.
@@ 4.1. Set up the Benchmark
Models. We compare VVO with both OCL classics and state-of-the-arts. SLATE (Singh et al., 2022a) uses a Transformer decoder for auto-regressive decoding, and it differs from VVO in a separate VAE encoder and naive quantizer; STEVE (Singh et al., 2022c) is SLATE’s video version. DINOSAUR (Seitzer et al., 2023) uses an MLP for mixture decoding, and it differs from VVO in no quantization in its reconstruction target. SlotDiffusion (Wu et al., 2023b) uses a conditional Diffusion model for diffusion decoding, and it differs from VVO in a separate VAE encoder and naive quantizer. General improvers GDR (Zhao et al., 2024a) and MSF (Zhao et al., 2024b) only support auto-regression and diffusion decoding. We skip outdated methods like IODINE (Greff et al., 2019), SA (Locatello et al., 2020) and ISA (Biza et al., 2023) due to their low accuracy. We also skip SAVi (Kipf et al., 2022) and SAVi++ (Elsayed et al., 2022) as their extra modalities are unfair to others.

### v6

@@ 1. Introduction
Objects can form highly diverse visual scenes through arrangements and combinations. But mainstream methods based on feature patches or a single feature vector disregard such compositionality. Inspired by human vision cognition (bar2004visual; cavanagh2011visual; palmeri2004visual), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects. Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (wu2022slotformer), as well as in interactions between visual modality and other modalities (zhang2024omgllava; wang2024omnidrive).
Existing OCL methods typically adopt an encoder-aggregator-decoder architecture (locatello2020slotattent). Firstly, the encoder transforms input image or video frame pixels into a dense feature map. Then, the aggregator sparsifies this feature map into feature vectors via Slot Attention (locatello2020slotattent) with initial slots as the query. Lastly, the decoder reconstructs the input in some form from these aggregated slots, to provide the self-supervised training signal.
OCL relies on pixel textures to discover objects. The early milestone (locatello2020slotattent) reconstructs input pixels as supervision, usually failing on realistic objects. Some (kipf2021savi; elsayed2022savipp) reconstruct optical flow or depth map to mitigate textural noises, at the cost of expensive annotations. Some (singh2021slate; singh2022steve) reconstruct input’s VAE (Variational Autoencoder) representation, whose super-pixels are codebook codes, thus suppressing pixel redundancy and facilitating aggregation from features into slots. Recent advances (seitzer2023dinosaur; wu2023slotdiffuz) use Vision Foundation Models (VFMs) (caron2021dino1; oquab2023dino2) to extract features with better object separability, boosting OCL significantly.
@@ 2. Related Work
OCL encoding. Early milestone methods like IODINE(greff2019iodine) and SA (locatello2020slotattent) use small naive CNNs (krizhevsky2012alexnet) as OCL encoder. Followups like SAVi (kipf2021savi), SAVi++ (elsayed2022savipp), SLATE (singh2021slate) and STEVE (singh2022steve) employ pretrained ResNets (he2016resnet), and fine-tune them on OCL datasets. State-of-the-arts like SlotDiffusion (wu2023slotdiffuz) and DINOSAUR (seitzer2023dinosaur) utilize VFMs like DINO (caron2021dino1) and DINO2 (oquab2023dino2) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly. SAM (kirillov2023sam) and SAM2 (ravi2024sam2) are also well recognized VFMs yet remain unexploited in OCL setting. Our VVO supports various VFMs for OCL encoding.
OCL aggregation. SlotAttention (locatello2020slotattent) is the footstone for mainstream OCL methods. Subsequent works like BO-QSA (jia2023boqsa), ISA (biza2023isa) and SysBind (singh2022sysbind) are all its variants, which are designed without changing the external interface. But considering their performance boosts, we only integrate BO-QSA by default.
OCL decoding. With SlotAttention as the aggregator, the decoder and its reconstruction target affect OCL performance the most, as it is the source of supervision. Mixture-based decoding, used in SAVi, SAVi++, DINOSAUR and VideoSAUR (zadaianchuk2024videosaur), decodes each slot’s spatial broadcast (watters2019spatialbroadcast) using naive CNNs or MLPs, and mixes them with corresponding weights into the reconstruction. Transformer-based decoding, used in SLATE, STEVE and SPOT (kakogeorgiou2024spot), reconstructs VAE representation of the input auto-regressively with slots as the condition. Diffusion-based decoding in LSD (jiang2023lsd) and SlotDiffusion drives slots to recover noise added to the input’s VAE representation. Our VVO supports all these types of OCL decoding.
VAE for OCL. Variational Autoencoders (VAEs), like dVAE (im2017dvae) in SLATE and VQ-VAE (van2017vqvae) in SlotDiffusion, are employed to produce reconstruction targets for OCL training. Since these VAEs are designed for image generation, some methods adapt them for OCL. Inspired by channel or weight grouping, GDR (zhao2024gdr) decomposes features into attributes and combine them to produce VAE representation as reconstruction targets to guide OCL better. MSF (zhao2024msf) firstly exploits the multi-scale idea in the OCL setting with VAE-specific designs. Based on recent advancement RVQ (yang2023rvq) and SimVQ (zhu2024simvq), we design our own VQ variant for OCL.
@@ 3.1. Unify OCL
where $\bm{\phi}_{\mathrm{e}}$ can be parameterized as pretrained VFMs, like DINO (caron2021dino1), DINO2 (oquab2023dino2), SAM (kirillov2023sam) and SAM2 (ravi2024sam2). As OCL relies on textures to separate objects, $\bm{\phi}_{\mathrm{e}}$ should handle complex textures of objects, making VFMs necessary here. We will explain it in Section 3.2.
@@
where $\bm{\phi}_{\mathrm{a}}$ can be parameterized as widely adopted SlotAttention (locatello2020slotattent) and its variants, which is some cross attention with $\bm{S}_{0}$ as queries and $\bm{Z}$ as keys and values. $\bm{M}$ is the binarized attention map thus intuitively reflects how well objects are represented by slots.
For video OCL, there is a recurrent module transitioning current slots $\bm{S}$ into new queries $\bm{S}_{0}$ for the next time step. Such module can be parameterized as a Transformer encoder block (vaswani2017transformer).
@@
where $\bm{\phi}_{\mathrm{q}}$ can be parameterized as some Vector Quantization (VQ) (im2017dvae; van2017vqvae). But we meticulously design our own VQ variant, as detailed in Section 3.2. $\bm{\phi}_{\mathrm{q}}$ is pretrained in VAE framework and is frozen afterwards, where the encoder is shared from the frozen OCL encoder $\bm{\phi}_{\mathrm{e}}$, and the decoder is a typical VAE decoder. We will explain why not use a separate typical VAE encoder in Section 3.2.
@@
Here $\bm{\phi}_{\mathrm{d}}$ can be parameterized as (i) a CNN or MLP for mixture decoding (kipf2021savi; seitzer2023dinosaur), where $\bm{Q}$ is destructed to height and width, and $\bm{S}$ is spatially broadcast (watters2019spatialbroadcast) into this shape and then decoded into components being mixed together; (ii) or a Transformer decoder for auto-regressive decoding (singh2021slate; kakogeorgiou2024spot), where $\bm{Q}$ is destructed with causal masking as the query and $\bm{S}$ is the key and value; (iii) or a conditional Diffusion model for diffusion decoding (wu2023slotdiffuz; jiang2023lsd), where $\bm{Q}$ is destructed with noise as the input and $\bm{S}$ is the condition.
Reconstructing $\bm{Q}$ using $\bm{S}$ drives $\bm{S}$ to aggregate as much object information as possible. Thus, a good reconstruction target, like in (zhao2024gdr; zhao2024msf), is very important. We will explain this in Section 3.2.
@@ 3.2. Utilize VFMs in OCL
We directly extract the feature map $\bm{Z}$ from the input $\bm{X}$ using VFMs as $\bm{\phi}_{\mathrm{e}}$, where DINO2 ViT and SAM2 encoder are chosen and experimented thoroughly. No extra position encoding is needed here like in SA (locatello2020slotattent) because these VFMs already contain the positional information required in $\bm{\phi}_{\mathrm{a}}$. Since $\bm{\phi}_{\mathrm{e}}$ is frozen, We further use a trainable linear layer to adjust $\bm{Z}$ slightly.
As shown in Figure 1, VFM representations have better objectness than non-VFMs, even under naive kMeans clustering [^1] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html. OCL aggregation is essentially a clustering whose initial centroids are trainable (jia2023boqsa). Thus, we expect $\bm{\phi}_{\mathrm{a}}$ to aggregate VFM’s $\bm{Z}$ into slots $\bm{S}$ better under queries $\bm{S}_{0}$. Previous methods like DINOSAUR (seitzer2023dinosaur) have already exploited this but without any reason.
@@
Our VQ’s codebook follows SimVQ (zhu2024simvq). We predefine $m=4096$ template features, i.e., a codebook $\bm{T}_{0}\in\mathbb{R}^{m\times c_{0}}$, which are randomly initialized and remain frozen. In vector quantization, we project $\bm{T}$ with a pre-trainable linear layer and match it with the adjusted $\bm{Z}$:
@@
where $\mathrm{softmax}(\cdot)$ is calculated along the channel dimension; $\bm{P}$ is the match probabilities; $\bm{I}\in\mathbb{R}^{h\times w}$ is the matched code indexes; $\mathrm{argmax}(\cdot)$ is calculated along the channel dimension; $\mathrm{index}(\cdot,\cdot)$ is operated along code number dimension. The typical STE (bengio2013ste) on $\bm{Q}$, needed in pre-training, can be skipped during OCL training.
@@
where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is Gumbel noise and $\tau$ is the temperature. Training-time annealing residual connection (zhao2024gdr; zhao2024msf) is added after Equation 10 to stabilize pre-training:
@@
where $\alpha$ is scheduled from 1 to 0 during pre-training using cosine-annealing. Besides typical losses of reconstruction, alignment and commitment (van2017vqvae), we regularize the adjusted $\bm{Z}$ to be normal:
@@
With all samples’ feature maps $\bm{Z}$ being represented with one codebook $\bm{T}$, the quantization $\bm{Q}$ naturally gains cross-sample consistency, helping the aggregation with queries $\bm{S}_{0}$, which are also shared across samples. Such tokenization is compatible with both regression and classification decoding. In contrast, methods like SLATE (singh2021slate) and SlotDiffusion (wu2023slotdiffuz) are faced with distribution gaps between $\bm{Q}$ and $\bm{Z}$, shown in Figure 1, due to separate VAE and OCL encoders. Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen OCL supervision.
@@ 3.3. Compare Architectures
- SLATE0 (the official version) (singh2021slate): (1) VFM features are NOT directly fed to the aggregator. (2) VFM features are shared discretized to scalar numbers and re-embedded into features to be learned latter. This loses much information of VFM features.
- SLATE (the improved version; adopted here) / STEVE (jia2023boqsa; singh2022steve; wu2023slotdiffuz): (1) Same as VVO. (2) Reconstruction targets are discretized from features of separate VAE encoding, not quantized from features shared from OCL encoding, causing optimization noises.
- SlotDiffusion (wu2023slotdiffuz) / LSD (jiang2023lsd): (1) Same as VVO. (2) Reconstruction targets are quantized from separate VAE encoding features, not sharing OCL encoding features, causing optimization noises.
- DINOSAUR (seitzer2023dinosaur) / VideoSAUR (zadaianchuk2024videosaur) / SPOT (kakogeorgiou2024spot): (1) Same as VVO. (2) Reconstruction targets are shared from VFM features of OCL encoding without quantization, causing optimization noises.
@@
- Mixture-based OCL decoders, e.g., CNN (locatello2020slotattent; kipf2021savi; elsayed2022savipp), MLP (seitzer2023dinosaur) and SlotMixer (zadaianchuk2024videosaur), are originally designed for continuous features, thus are compatible with our shared quantized VFM features.
- Auto-regressive OCL decoders, e.g., the Transformer decoder (singh2021slate; singh2022steve) and Transformer9 (kakogeorgiou2024spot), are designed for discretized features, while also showing applicability to continuous features (seitzer2023dinosaur; kakogeorgiou2024spot). Thus they are applicable to our shared quantized VFM features.
- Diffusion-based OCl decoders, e.g., conditional Diffusion (wu2023slotdiffuz; jiang2023lsd), work on low-dimension features, necessitating our shared quantization on the continuous high-dimensional VFMs features.
@@ 4.1. Set up the Benchmark
Models. We compare VVO with both OCL classics and state-of-the-arts. SLATE (singh2021slate) uses a Transformer decoder for auto-regressive decoding, and it differs from VVO in a separate VAE encoder and naive quantizer; STEVE (singh2022steve) is SLATE’s video version. DINOSAUR (seitzer2023dinosaur) uses an MLP for mixture decoding, and it differs from VVO in no quantization in its reconstruction target. SlotDiffusion (wu2023slotdiffuz) uses a conditional Diffusion model for diffusion decoding, and it differs from VVO in a separate VAE encoder and naive quantizer. General improvers GDR (zhao2024gdr) and MSF (zhao2024msf) only support auto-regression and diffusion decoding. We skip outdated methods like IODINE (greff2019iodine), SA (locatello2020slotattent) and ISA (biza2023isa) due to their low accuracy. We also skip SAVi (kipf2021savi) and SAVi++ (elsayed2022savipp) as their extra modalities are unfair to others.